OpenStudy (anonymous):

A rectangular pen is made up of 150 feet of fencing. Can someone please draw a sketch of th pen and label the dimensions. Express the area of the enclosure as a function of the width.

OpenStudy (anonymous):

 \[A=LW\] \[2W+2L=150\] \[2W=150-2L\] \[L=75-W\] \[A(W)=W(75-W)\]

OpenStudy (anonymous):

are you sure? and is this how you do this?

OpenStudy (anonymous):

yes i am pretty sure. is there more to the question? you can also write \[A=75W-W^2\] if you like

OpenStudy (anonymous):

yes theres more..it asks about the window

OpenStudy (anonymous):

xmax, xmin, xscl, ymin, ymax yscl..

OpenStudy (anonymous):

what window? are you supposed to graph this on a calculator?

OpenStudy (anonymous):

that looks like calculator settings for graphing.

OpenStudy (anonymous):

i lost my calculator...the window is just the xmax, xmin, etc

OpenStudy (anonymous):

i don't know what xmax, xmin etc is if it is not the window on a calculator.

OpenStudy (anonymous):

its is the window on the calculator....but I couldnt find mine..

OpenStudy (anonymous):

thnx for helping